# CoordAR: One-Reference 6D Pose Estimation of Novel Objects via Autoregressive Coordinate Map Generation

### v2

@@ Introduction
  • We are the first to introduce autoregressive coordinate map generation for 6D pose estimation of novel objects. We further demonstrate the superiority of autoregressive generation compared to parallel real-valued regression.
@@
The YCB-V (xiang2017posecnn) dataset contains numerous occluded scenes, including cases where even the first reference view is occluded. As displayed in rows 4 to 7 of Fig. 4, our method exhibits robustness when the query and reference are occluded. We also observe that our method performs well when the object frame is ill-defined (see row 6), suggesting that our method is independent of the definition of the canonical object frame. As demonstrated in Tab. 2, we observe a slightly lower ADD AUC compared to One2Any (liu2025one2any). Note that YCB-V contains texture-rich food containers that are geometrically symmetric but have different texture-influenced symmetry definitions during evaluation, such as the tomato_soup_can. In such cases, ADD AUC may penalize predictions that are geometrically correct but differ in texture alignment (see the appendix for more details). Nevertheless, our method achieves a significantly higher ADD-S AUC than existing single-view methods, which is important for applications that emphasize geometric perception.
@@
Our model supports trade-offs between accuracy and computational efficiency by adjusting the number of generation steps. Inference speed comparisons, along with their ADD(-S) on the LINEMOD dataset, are provided in Tab. 5. Impressively, our model can achieve near real-time speed (0.10 seconds per frame) with even a single step while maintaining comparable accuracy.
@@ Acknowledgments
This work was supported in part by National Key R&D Program of China under Grant 2022YFB3903801, in part by the National of Science Foundation of China under Grant 62073214, and in part by the Corporate Research Center, State Key Laboratory of High-end Heavy-load Robots, Midea Group (3D Robot Vision Project). We thank Kun Wang, Li Shen, Yuhui Ni, and Yikun Zeng for providing the reproduced results of baseline methods for qualitative comparison.
@@
In this supplementary, we provide detailed information about data preparation, method implementation, and experimental results.

### v3

@@ Introduction
  • To the best of our knowledge, we are the first to introduce autoregressive coordinate map generation for 6D pose estimation of novel objects. We further demonstrate the superiority of autoregressive generation compared to parallel real-valued regression.
@@
The YCB-V (xiang2017posecnn) dataset contains numerous occluded scenes, including cases where even the first reference view is occluded. As displayed in rows 4 to 7 of Fig. 4, our method exhibits robustness when the query and reference are occluded. We also observe that our method performs well when the object frame is ill-defined (see row 6), suggesting that our method is independent of the definition of the canonical object frame. While we observe a slightly lower ADD AUC in Tab. 2 compared to One2Any (liu2025one2any), our method achieves a significantly higher ADD-S AUC than existing single-view methods. We attribute this discrepancy to a bias in the YCB-V evaluation protocol. Note that YCB-V contains texture-rich food containers that are geometrically symmetric but have different texture-based symmetry definitions during evaluation, such as the tomato_soup_can. In such cases, ADD AUC can penalize predictions that are geometrically correct but differ in texture alignment, even though they achieve strong performance under the ADD-S metric (see the full YCB-V results in the appendix for details). Interestingly, when switching to the first frame of each scene as the reference, our method outperforms One2Any in both ADD AUC and ADD-S AUC, as demonstrated by the pose tracking results reported in the appendix.
@@
Our model supports trade-offs between accuracy and computational efficiency by adjusting the number of generation steps. Inference speed comparisons, along with their corresponding ADD(-S) on the LINEMOD dataset, are provided in Tab. 5. Impressively, our model can achieve near real-time speed (0.10 seconds per frame) with even a single step while maintaining comparable accuracy.
@@ Acknowledgments
This work was supported in part by the National Key R&D Program of China under Grant 2022YFB3903801, in part by the National Science Foundation of China under Grant 62073214, and in part by the Corporate Research Center, State Key Laboratory of High-end Heavy-load Robots, Midea Group (3D Robot Vision Project). We thank Kun Wang, Li Shen, Yuhui Ni, and Yikun Zeng for providing the reproduced results of baseline methods for qualitative comparison.
@@
In this appendix, we provide detailed information about data preparation, method implementation, and experimental results.